OpenStudy (watchmath):

Compute \[\lim_{x\to\-\infty}\sqrt{x^2+x}\sin(1/x)\]

OpenStudy (anonymous):

1

OpenStudy (watchmath):

Let me make it better \[\lim_{x\to -\infty}\sin(1/x)\sqrt{x^2+x}\]

OpenStudy (watchmath):

no :)

OpenStudy (anonymous):

oh! -1

OpenStudy (watchmath):

yup :D

OpenStudy (anonymous):

ok , lemme betta get some sleep :) and wer r u from watchman?

OpenStudy (watchmath):

I will say that I am in US right now :D

OpenStudy (anonymous):

I got 0 times infinity; how did you get to -1?

OpenStudy (watchmath):

write it as \[\frac{\sin(1/x)}{1/x}\cdot \frac{\sqrt{x^2+x}}{x}\]

OpenStudy (anonymous):

how did you make that manipulation?

OpenStudy (anonymous):

Oh, I see, 1/x times x is equal one, no change; but where does the negative come from?